# A Comprehensive Multi-Vocal Empirical Study of ML Cloud Service Misuses

### v1

@@ Abstract.
Machine Learning (ML) models are widely used across various domains, including medical diagnostics and autonomous driving. To support this growth, cloud providers offer ML services to ease the integration of ML components in software systems. The evolving business requirements and the popularity of ML services have led practitioners of all skill levels to implement, and maintain ML service-based systems. However, they may not always adhere to optimal design and usage practices for ML cloud services, resulting in common misuse which could significantly degrade the quality of ML service-based systems and adversely affect their maintenance and evolution. Though much research has been conducted on ML service misuse, a consistent terminology and specification for these misuses remain absent. We therefore conduct in this paper a comprehensive, multi-vocal empirical study exploring the prevalence of ML cloud service misuses in practice. We propose a catalog of 20 ML cloud service misuses, most of which have not been studied in prior research. To achieve this, we conducted a) a systematic literature review of studies on ML misuses, b) a gray literature review of the official documentation provided by major cloud providers, c) an empirical analysis of a curated set of 377 ML service-based systems on GitHub, and d) a survey with 50 ML practitioners. Our results show that ML service misuses are common in both open-source projects and industry, often stemming from a lack of understanding of service capabilities, and insufficient documentation. This emphasizes the importance of ongoing education in best practices for ML services, which is the focus of this paper, while also highlighting the need for tools to automatically detect and refactor ML misuses.
$c$s: ML Service, Misuse, Software Quality
@@ 1. Introduction
Major breakthroughs in Artificial Intelligence (AI) in general, and Machine Learning (ML) in particular, have brought huge success to many application domains such as image recognition, security, and autonomous driving (grigorescu2020survey, ; dong2021survey, ).
@@
To support this technological hype, several major cloud providers like Amazon, Google, and Microsoft, have been pivotal in providing ML practitioners with various ML cloud infrastructure capabilities and specialized services necessary for performing ML-related tasks (wan2021machine, ). The availability of ML cloud services, which do not require designing or building ML models from scratch, has significantly facilitated the adoption and integration of ML in software systems. Existing ML cloud services cover a wide spectrum of services, from extreme simplicity and complete customizability, for building ML service-based systems (google_cloud_automl, ; ibm_watson, ; ibm_data_platform, ). For instance, some ML cloud services operate as black-box systems, not disclosing the ML classification models they employ, while others offer support throughout ML development stages, from data preprocessing to the selection of ML classifiers, parameter tuning, and model deployment (azure_ml_designer, ). The growing demand for business solutions and the rising popularity of ML services have driven practitioners of diverse skill levels to adopt them to speed up the development, maintenance and evolution of ML-based systems (wan2021machine, ). However, this rapid pace may cause inconsistencies in adhering to best practices for designing and utilizing ML services (wan2021machine, ). As a result, this can lead to service misuse, which can negatively impact the quality of ML service-based systems and pose challenges in their maintenance and evolution (washizaki2022software, ). More specifically, ML service misuse may introduce severe bugs that can affect the correctness, performance, and cost-efficiency of ML service-based systems (wan2021machine, ).
@@
ML service misuse is defined as poor practices in the use of services within the ML development pipeline, including data collection and preprocessing, training, testing, deployment, serving, and monitoring. These misuses are also perceived as a violation of (implicit) service usage constraints in end systems, potentially leading to inefficiencies or unintended consequences on end systems. A recent study reported that 69% of projects on GitHub using Amazon and Google ML cloud services contain service misuses that directly impact the system’s functionality and service costs (wan2021machine, ). For example, failing to specify a stopping criterion during the training of models through an ML service can increase latency as well as cloud usage costs, as training might continue unnecessarily long without significant performance improvements (intro, ). Another example is the availability of synchronous and asynchronous versions of the same ML task offered by several ML cloud providers. Asynchronous services support larger inputs and longer processing times than synchronous services. However, invoking asynchronous ML services in a synchronous manner can drastically impact the performance of an ML service-based system by increasing its latency due to blocked service invocations. Moreover, ML cloud providers typically establish rate limits for ML services, setting the maximum permissible requests within a certain time frame, which helps in regulating ML service utilization and prevents system overloads. However, exceeding such limits without proper handling can result in critical issues, such as increased latency and instability in ML service-based systems. Therefore, it is important to understand ML service misuse in software systems to promote better improvement of software quality and maintain higher levels of correctness, performance, and cost-efficiency.
@@
Unlike ML or service antipatterns, which have well-established definitions in prior research (washizaki2022software, ; bogner2021characterizing, ), ML service misuses still lack precise specifications and a comprehensive understanding in practical contexts. Additionally, existing research on ML service misuses is limited, with the highest number of misuses reported by a single study being eight. (wan2021machine, ). Also, the terminology used by prior research is often inconsistent and have vague definitions, which impedes practitioners from fully understanding and addressing any relevant issues. Recognizing these misuses can help increase awareness of common pitfalls when developing ML service-based systems, thereby potentially improving their overall quality.
@@
Therefore, in this paper, we conduct a comprehensive, multi-vocal empirical study on common misuses in ML cloud services. We propose a catalog of 20 ML service misuses, drawing from a systematic literature review, a gray literature review, and an analysis of ML service-based systems on GitHub. To validate our catalog, we conducted an online survey with 50 ML practitioners, in which we aimed to not only confirm the identified misuses but also to explore their prevalence in industry, understand the underlying causes, and gain insights from practitioners regarding potential mitigation strategies. Our empirical analyses revealed a strong agreement among practitioners on the validity and relevance of our catalog. The results show that ML service misuses are widely spread in industry, with “Improper handling of ML API limits” being the most frequent misuse. Practitioners indicated that these misuses typically arise from a lack of understanding of ML service capabilities, insufficient cloud documentation, and limited awareness of best practices for ML services. To address these challenges, practitioners emphasized the importance of regular performance reviews to maintain efficient ML service-based systems. They also advocated for thorough validation of data quality, combined with continuous education on best practices, to ensure the correct use of ML services.
@@
  • We conducted a systematic review of the research literature related to ML systems, cloud services, and (anti)patterns, which resulted in 39 relevant papers, revealing a limited number of ML service misuse scenarios.
  • We conducted a gray literature review of the official documentation provided by major ML cloud providers, which helped identify a more comprehensive set of ML service misuses that have not been studied before.
  • We performed an in-depth empirical analysis of a curated set of 87 GitHub repositories and employed a mixture of manual and static code analysis to identify ML service misuses, resulting in a validated dataset of such misuses.
@@
The rest of this paper is structured as follows. Section 2 presents the related work. Section 3 describes the study design. The catalog proposal and the analysis of survey results are presented in Section 4. Section 5 discuss the implications of our findings. Section 6 discusses the potential validity threats to our findings. Finally, Section 7 concludes our work and suggests possible future research.
@@ 2. Related Work
The use and misuse of ML cloud services remain largely unexplored, with a few exceptions. Wan et al. (wan2021machine, ) analyzed 360 open-source software systems that use Google and AWS ML APIs and identified eight ML API misuses that negatively impact the quality of these systems. However, the authors only analyzed ML cloud APIs that give access to pre-trained models. Unlike our work, ML services misuse related to data preprocessing, model training, and model deployment were not covered. In addition, some misuses were generic in the sense that they apply to any cloud API, not exclusively those related to ML. As a result, in our paper, we considered only four misuses from their study. Moreover, Wei et al. (wei2024demystifying, ) conducted a study on Deep Learning (DL) API misuse in Python libraries such as TensorFlow and PyTorch, exploring their characteristics and proposing an approach to detect them. However, their work was not specific to ML service misuses.
Other research studies addressed the specification of ML (anti)patterns (masuda2018survey, ; washizaki2019studying, ; bogner2021characterizing, ). In particular, Washizaki et al. (washizaki2019studying, ) conducted a literature review and a survey with developers to collect, classify, and discuss (anti)patterns for ML-based systems. The authors found that developers have little knowledge of ML design patterns that could aid in developing their software systems. Bogner et al. (bogner2021characterizing, ) conducted a literature review to study technical debts in ML-based systems and antipatterns (cabral2024investigating, ). The authors argued that there is still no comprehensive conceptual overview of technical debts in ML-based systems, despite existing research efforts. They compiled a set of 72 ML antipatterns aimed at avoiding technical debt in the literature. However, they did not specifically study ML service antipatterns.
@@
With respect to the software quality of ML-based systems, there has been little research in the literature that highlighted the need for adapted quality models (wang2024quality, ; masuda2018survey, ; lewis2021software, ; kuwajima2020engineering, ). For example, Masuda et al. (masuda2018survey, ) summarized software quality methods and techniques for ML-based systems. Kuwajima et al. (kuwajima2020engineering, ) studied the suitability of a traditional software quality model for ML systems and provided insights into the major differences between ML systems and conventional software systems. Cardozo et al. (cardozo2023prevalence, ) analyzed the code quality of reinforcement learning projects by utilizing software metrics related to the definition, access, and interaction of program entities. These metrics served as a proxy of code quality, aiding in the detection of code smells. We leveraged the findings in these studies to specifically target ML service misuses that can negatively impact the quality of software systems. Identifying such misuses would greatly help increase the awareness of common pitfalls when building ML service-based systems and improve their overall quality.
3. Study Design
@@
Figure 1 depicts our study design to build our catalog of ML service misuses. It consists of four steps, which we will describe in details in the following section.
Figure 1. Overview of the Study Design
3.1. Systematic Literature Review
Following the guidelines of Kitchenham et al. (kitchenham2004procedures, ) to conduct systematic literature reviews, we first collected research papers based on a search query. First, we identified key terms related to ML systems, then identified a corresponding set of synonyms that are connected. We thus defined the following search query:
(Machine Learning OR ML OR Artificial Intelligence OR AI OR Deep Learning OR DL) AND
(Misuse OR Antipattern OR Bad Practice OR Pitfall OR Best Practice OR Pattern OR Software Quality OR
Quality Assurance OR Performance) AND (Cloud OR Azure OR Google OR AWS OR Service OR API OR SDK)

### v2

@@ Abstract.
Machine Learning (ML) models are widely adopted across many application domains. To support this trend, major cloud providers offer ML cloud services that eliminate the need to build models from scratch. These ML services empower practitioners with varying expertise to rapidly embed ML capabilities into software systems. However, practitioners often overlook best practices and optimal design when using cloud-based services. This results in recurring misuses that can compromise system quality and may hinder long-term maintenance and evolution. Though prior work has studied some misuse cases, the field lacks consistent terminology and clear specifications of ML service misuses. In this paper, we address three research questions (RQs) investigating (1) the types of ML service misuses have been identified in research literature, gray literature, and open-source software systems, (2) the state of the practice of ML cloud service misuses in industry, and (3) the prevalence of ML service misuses in open-source projects in comparison with what is observed in industry practice. To address these RQs, we conducted a comprehensive multi-vocal empirical study investigating the prevalence of ML cloud service misuses in practical contexts. Overall, our analyses incorporated a diverse range of sources, including academic research, official documentation from major ML cloud providers, and 377 GitHub projects using ML services, and were complemented by a survey of 50 ML practitioners from industry. As a result, we developed a catalog of 20 ML service misuses, cross-validated through evidence from both open-source projects and industry feedback. We observe that our identified misuses are widely prevalent in both open-source projects and industry, often due to a lack of understanding of service capabilities, inadequate documentation, and poor awareness of best practices. This highlights the importance of ongoing education on best practices for ML services and highlights the need for tools to automatically detect and refactor ML service misuses.
ML Could Services, Misuse, Software Quality
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Software and its engineering Software design engineeringccs: Computing methodologies Machine learningccs: Software and its engineering Software system structuresccs: Information systems Cloud computing
@@ 1. Introduction
Major breakthroughs in Artificial Intelligence (AI) in general, and Machine Learning (ML) in particular, have brought huge success to many application domains such as image recognition, security, and autonomous driving (grigorescu2020survey, ; dong2021survey, ). To support this technological hype, several major cloud providers, such as Amazon, Google, and Microsoft, have been pivotal in providing ML practitioners with various ML cloud infrastructure capabilities and specialized services that support key ML development tasks (wan2021machine, ). The availability of ML cloud services has greatly facilitated the adoption and integration of ML in software systems. These services provide not only pre-trained models and ready-to-use APIs but also powerful Platform-as-a-Service (PaaS) solutions for developing ML-based systems. Vertex AI (Google_VertexAI_Docs, ) and SageMaker (AWS_SageMaker_Train, ) are examples of platforms that allow developers to build, train, and deploy custom ML models more efficiently by offering managed infrastructure, specialized tools, and integrated workflows. This support accelerates the development of ML-based systems and simplifies the integration of ML components into software systems (DBLP:journals/corr/abs-1904-12054, ; DBLP:journals/corr/abs-2012-08483, ; DBLP:journals/corr/abs-2012-08489, ).
@@
Existing ML cloud services cover a wide spectrum of services, from extreme simplicity and complete customizability, to build ML service-based systems (google_cloud_automl, ; ibm_watson, ; ibm_data_platform, ). For instance, some ML cloud services operate as black-box systems (i.e., not disclosing the ML classification models they employ) while others offer support throughout ML development stages, from data preprocessing to the selection of ML classifiers, parameter tuning, and model deployment (azure_ml_designer, ). The growing demand for ML-driven business solutions has led practitioners of varying expertise to increasingly adopt cloud-based ML services to accelerate the development, maintenance, and evolution of ML systems (wan2021machine, ). However, this rapid pace often comes at the cost of adhering to best practices for designing and utilizing ML services (wan2021machine, ), resulting in recurring misuses that can degrade system quality, pose challenges in their maintenance and evolution, introduce bugs, and impact performance and cost-efficiency (washizaki2022software, ).
@@
ML service misuse is defined as poor practices in the use of services within the ML development pipeline, including data collection and preprocessing, training, testing, deployment, serving, and monitoring. These misuses are also perceived as a violation of (implicit) service usage constraints in end systems, potentially leading to inefficiencies or unintended consequences on end systems. A recent study reported that 69% of GitHub projects using Amazon and Google ML cloud services contain service misuses that directly impact system functionality and service costs (wan2021machine, ). As an example, failing to specify a stopping criterion during model training through an ML service can increase latency and cloud usage costs, as training might continue unnecessarily long without significant performance improvements (intro, ). Another example is the availability of synchronous and asynchronous versions of the same ML task offered by several ML cloud providers. Figure 1 shows an example of such a misuse with Azure cognitive services for real-time sentiment analysis of short text messages. As illustrated in the figure, asynchronous services support larger inputs and longer processing times than synchronous services. These two API types differ in their request handling and implications for system performance. The synchronous API allows immediate processing of user requests and returns sentiment results with minimal latency, offering a responsive and cost-efficient user experience. In contrast, the asynchronous API requires the submission of a job and the repeated polling of the results, leading to longer processing times, higher resource usage, and poorer user experience when misused as a synchronous service.
Figure 1. Synchronous and Asynchronous API: Correct Use vs. Misuse
Also, ML cloud service providers usually set rate limits for their services. These limits define the maximum allowable requests within a specific time frame, helping to regulate ML service usage and prevent system overloads. However, exceeding such limits without proper handling can result in critical issues, such as increased latency and instability in ML service-based systems. Therefore, understanding ML service misuse in software systems is important for improving software quality, ensuring correctness, improving performance, and achieving cost efficiency.
@@
Unlike ML or service antipatterns, which have well-established definitions in prior research (washizaki2022software, ; bogner2021characterizing, ), ML service misuses still lack precise specifications and comprehensive understanding in practical contexts. In addition, existing research on ML service misuse is limited, with the highest number of misuses reported by a single study being eight (wan2021machine, ). Besides, the terminology used by prior research is often inconsistent and has vague definitions, which prevents practitioners from fully understanding and addressing any relevant issues. Recognizing these misuses can help increase awareness of common pitfalls when developing ML service-based systems, thereby potentially improving their overall quality. Therefore, in this paper, we conduct a comprehensive, multi-vocal empirical study on common misuses in ML cloud services. Specifically, we address three research questions (RQs) that investigate: 1) the types of ML service misuses have been identified in research literature, gray literature, and open-source software systems, (2) the state of the practice of ML cloud service misuses in industry through an online survey with 50 ML practitioners, and (3) the prevalence of ML service misuses in open-source projects in comparison with what is observed in industry practice. We developed a comprehensive catalog of 20 misuses, categorized by the ML development pipeline stage. We explored the prevalence of our catalog in the industry to understand the underlying causes and gain insights from practitioners on potential mitigation strategies.
@@
Our empirical analyses revealed a strong agreement among practitioners on the validity and relevance of our catalog. The results show that ML service misuses are widely spread in industry, with “Improper handling of ML API limits” being the most frequent misuse. Practitioners indicated that these misuses typically arise from a lack of understanding of ML service capabilities, insufficient cloud documentation, and limited awareness of best practices for ML services. To address these challenges, practitioners emphasized the importance of regular performance reviews to maintain efficient ML service-based systems. They also advocated for a thorough validation of data quality, combined with continuous education on best practices, to ensure the correct use of ML services.
@@
  • We conducted a thorough review of the research literature related to ML systems, cloud services, and (anti)patterns, which resulted in 31 relevant papers, revealing a limited number of ML service misuse scenarios.
  • We conducted a gray literature review to identify a more comprehensive set of ML service misuses that have not been previously studied in the context of ML service-based systems.
  • We conducted an in-depth empirical analysis of a curated set of 377 GitHub repositories, employing a combination of manual and static code analysis to identify misuses of ML services, resulting in a validated dataset of such misuses.
@@
The remainder of this paper is structured as follows. Section 2 presents the related work. Section 3 highlights the motivation of our work. Section 4 describes our multi-vocal study. The catalog proposal and the analysis of the survey results are presented in Section 5. Section LABEL:Sec:discussion discusses the implications of our findings. Section LABEL:Sec:threats discusses the potential validity threats to our findings. Finally, Section LABEL:sec:ccl concludes our work and suggests possible future research.
@@ 2. Related Work
This section provides an overview of the key literature related to our study. We discuss previous work in three areas: (1) ML API misuses, (2) ML technical debts, since certain misuses can introduce such debts, and (3) ML quality models, as certain ML misuses can have an impact on software quality attributes.
2.1. ML API misuses
@@
The use and misuse of ML cloud services remain largely underexplored, with a few exceptions. Wan et al. (wan2021machine, ) analyzed 360 open-source software systems that use Google and AWS ML APIs and identified eight ML API misuses that negatively impact the quality of these systems. However, the authors only analyzed ML cloud APIs that give access to pre-trained models. Unlike our work, ML service misuses related to data preprocessing, model training, and model deployment were not covered. In addition, some misuses were generic in the sense that they correspond to any cloud API, not exclusive to ML APIs. As a result, only four misuses from their study were considered in our study. Another study was conducted by Wei et al. (wei2024demystifying, ), in which they covered Deep Learning (DL) API misuse in Python libraries, such as TensorFlow and PyTorch, exploring their characteristics and proposing an approach to detect them. However, their work did not specifically address misuse issues related to ML cloud services. Other research studies addressed the specification of ML (anti)patterns (masuda2018survey, ; washizaki2019studying, ; bogner2021characterizing, ). In particular, Washizaki et al. (washizaki2019studying, ) conducted a literature review and a survey with developers to collect, classify, and discuss (anti)patterns for ML-based systems. The authors found that developers have little knowledge of ML design patterns that could aid in developing their software systems. Zhang et al. (zhang2022code, ) explored the gap in software engineering practices for machine learning by introducing a catalog of 22 ML-specific code smells. They systematically gathered these smells from research literature, gray literature, commits, and Q&A platforms, and classified them by affected pipeline stage, long-term risks, and possible fixes to support better code maintainability in ML systems. In contrast, our work focuses specifically on ML misuses in cloud-based environments, which introduce unique challenges not addressed by Zhang et al.’s taxonomy. Furthermore, we complement our catalog with a practitioner survey to assess the level of agreement on these misuses and their perceived frequency in real-world cloud-based ML systems.
2.2. ML technical debts
Bogner et al. (bogner2021characterizing, ) conducted a literature review to study technical debts in ML-based systems and antipatterns. The authors argued that there is still no comprehensive conceptual overview of technical debts in ML-based systems, despite existing research efforts. They compiled a set of 72 ML antipatterns aimed at avoiding technical debt in the literature. However, they did not specifically study ML service antipatterns. Sculley et al. (sculley2015hidden, ) discussed the concept of ML technical debt by highlighting the long-term maintenance challenges and risks associated with deploying ML systems in real-world environments. They identified several ML-specific risk factors that contribute to system-level antipatterns, emphasizing the hidden complexities that arise when ML components are integrated into production systems. Moreover, Recupito et al. (recupito2024technical, ) continued the investigation of ML technical debt by focusing on the practical challenges faced by ML practitioners and identifying recurring technical issues that contribute to long-term software quality degradation. While these studies provide valuable insights into ML-related technical debt, we aim in our study to provide a comprehensive catalog of ML service misuses, which can introduce technical debt by embedding hidden complexities and long-term maintenance challenges into the system. By investigating the prevalence of these misuses in both open-source and industrial projects, we seek to understand their role in accelerating the accumulation of technical debt in ML services. In addition, we will analyze the underlying causes of these misuses and evaluate possible mitigation strategies from an industrial perspective, offering actionable guidance to potentially minimize technical debt in ML systems.
2.3. Quality models of ML-based systems
@@
With respect to the software quality of ML-based systems, there has been little research in the literature addressing the need for adapted quality models (wang2024quality, ; masuda2018survey, ; lewis2021software, ; kuwajima2020engineering, ). For example, Masuda et al. (masuda2018survey, ) summarized software quality methods and techniques for ML-based systems. Kuwajima et al. (kuwajima2020engineering, ) studied the suitability of a traditional software quality model for ML systems and provided insights into the key differences between ML systems and conventional software systems. Cardozo et al. (cardozo2023prevalence, ) analyzed the code quality of reinforcement learning projects by utilizing software metrics related to the definition, access, and interaction of program entities. These metrics served as a proxy of code quality, aiding in the detection of code smells. Cabral et al. (cabral2024investigating, ) investigated the impact of applying SOLID design principles to machine learning code, highlighting that adherence to these principles can significantly improve code understanding among data scientists. Their study suggests that established software engineering practices can improve maintainability and readability in ML projects, even when developed by practitioners of diverse educational backgrounds, emphasizing the importance of promoting design principles within the data science community. We leveraged the findings in these studies to specifically target ML service misuses that can negatively impact the quality of software systems. Identifying such misuses would greatly help increase awareness of common pitfalls when building ML service-based systems and improve their overall quality.
Summary on related work. Despite the increasing focus on antipatterns, technical debt, and quality concerns in ML-based systems, there remains a gap in identifying and characterizing misuses specific to cloud-based ML services across the ML lifecycle. Existing studies often focus on general ML design issues, local code-level antipatterns, or traditional ML frameworks, paying little attention to the unique practices and risks introduced by modern cloud-based ML service platforms. Our study addresses this gap by conducting a multi-vocal study on ML cloud services misuses, spanning from data processing to deployment and monitoring. Through a combination of research literature review, gray literature analysis, a practitioner survey, and examination of real-world projects, our aim is to raise awareness of recurring misuse patterns and support the adoption of ML services best practices in software systems.
3. Motivation
Major cloud providers now offer a broad range of ML cloud services, enabling developers at all skill levels to integrate advanced ML capabilities into software systems. This accessibility has greatly accelerated the adoption and integration of ML technologies. However, with increasing business demands and the growing prevalence of ML services, developers are building, maintaining, and evolving ML-based systems at a rapid pace, often without adhering to established best practices. Such misuse of ML cloud services can significantly degrade system quality, making them more difficult to maintain and evolve (wan2021machine, ). While prior research has mainly examined (anti)patterns, quality concerns, and technical debts in ML-based systems, there has been no systematic identification and characterization of misuses specific to ML cloud services. Although prior work has studied some misuse cases, there is still a lack of consistent terminology and clear specifications of ML service misuses in both academia and industry. In this paper, we aim to address this gap through a comprehensive multi-vocal empirical study on ML cloud service misuses. Our investigation combines (1) a review of both research and gray literature, (2) an empirical analysis of open-source GitHub projects, and (3) a survey of industrial practitioners to gain a deeper understanding of the current state of ML cloud service misuses. Our multi-vocal study is guided by the following research questions (RQs):
  • RQ1. What types of ML service misuses have been identified in research literature, gray literature, and open-source software systems? We aim to identify and categorize ML cloud service misuses described in both research and gray literature, as well as those observed in open-source software projects. Our objective is to build a comprehensive catalog that details the definition, characteristics, and concrete implementation examples of each misuse.
  • RQ2. What is the state of the practice of ML cloud service misuses in industry? Our goal is to explore practitioners’ perceptions of ML cloud service misuses that we identify in RQ1. Our goal is to investigate their agreement on such misuses as well as their prevalence in industry. Besides, we aim to investigate the causes practitioners attribute to such misuses and the mitigation strategies they employ.
  • RQ3. How comparable is the prevalence of ML service misuses in open-source projects to what is observed in industry practice? Our objective is to compare the prevalence of ML cloud service misuses in open-source systems with their occurrence in industry practice, to understand differences and similarities in misuse patterns across these environments.
4. Methodology
@@
Figure 2 depicts the design of our multi-vocal study to build and analyze our catalog of ML service misuses. The study consists of five main steps, which we describe in detail in the following subsections.
Figure 2. Overview of our methodology
4.1. Research Literature Review (RQ1)
To address our first research question, we conducted a structured review of the literature on the use and quality of machine learning systems in cloud-based environments. To ensure comprehensiveness and avoid missing relevant studies, we followed the approach guided by established principles and the guidelines of Kitchenham et al. (kitchenham2004procedures, ). Figure 3 depicts our process for paper selection. Initially, we defined a set of keywords related to our study, including machine Learning, misuse, and cloud services. Then, we identified a corresponding set of synonyms for each keyword and formulated the following search query:
s
("machine learning service" OR "ML service" OR "AI service" OR "deep learning service" OR "ML API" OR "AI API" OR "machine learning as a service" OR "MLaaS") AND (misuse OR smell OR "anti-pattern" OR "bad practice" OR "pitfall" OR "fault" OR "error" OR "Software Quality" OR "Quality Assurance") AND ("cloud" OR "Azure" OR "AWS" OR "Google Cloud")
Figure 3. Papers selection process of our research literature review